Hi! Please help solve the following problem:
Solve for x: 4x2+2x-7 = 13x+4
p.s. the 2 in 4x2 is an exponent, not a number.
I tried factoring and isolating the variable x to one side of the equation so that I could solve.
Thanks

Answer by checkley71(8403)

4X^2+2X-7=13X+4
4X^2+2X-13X-7-4=0
4X^2-11X-11 SOLVING USING THE QUADRATIC EQUATION X=(-B+-SQRT[B^24AC])/2A WE GET
X=(11+-SQRT[121-4*4*-11])/2*4
X=(11+-SQRT[121+176])/8
X=(11+-SQRT[297)/8
X=(11+-17.23)/8
X=(11+17.23)/8
X=28.23/8
X=3.528 SOLUTION
X=(11-17.23)/8
X=-6.23/8
X=-.778 SOLUTION